OpenStudy (anonymous):

which of the following is the equation of the exponential function y=Ce^kx that passes through (0,1) and (1,1/e) a)y=e^kx b)y=ce^x c)y=e^x d)y=e^-x e)none

OpenStudy (anonymous):

why u dont just replace (0,1) and (1,1/e) into the functions and see what match¡?

OpenStudy (anonymous):

i tried it already but i did not make sense

OpenStudy (anonymous):

really? the (d) dont match?

OpenStudy (agent0smith):

plug in (0, 1) \[\Large 1=Ce^{k \times 0}\]giving \[\Large 1 = C\]

OpenStudy (agent0smith):

Now plug in (1, 1/e) \[\Large \frac{ 1 }{ e }=e^{k \times 1}\]\[\Large \frac{ 1 }{ e }=e^{k}\] either take logs of both sides, or do this\[\Large e^{-1} = e^k\]

OpenStudy (anonymous):

so then i put them together

OpenStudy (agent0smith):

Yes, now you have both C and k.

OpenStudy (anonymous):

for the last problem you were doing was the answer k=-1

OpenStudy (agent0smith):

Yep :)

OpenStudy (anonymous):

oh ok so i first plug in the values then i solve for the values of c and k i think i get it now

OpenStudy (agent0smith):

Which equation did you get?

OpenStudy (anonymous):

i got d y=e^-x

OpenStudy (agent0smith):

:)